OpenStudy (anonymous):

If (89)p = 818, what is the value of p? 2 9 10 18

Nnesha (nnesha):

what do you think ? try it first!

OpenStudy (anonymous):

(8^9)^p = 8^18

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9!

OpenStudy (freckles):

\[8^{9^p}=8^{18} \text{ or } (8^9)^p=8^{18}\]

OpenStudy (anonymous):

cause u add the exponents

Nnesha (nnesha):

yes right \ remember the exponent rule \[\huge\rm (x^m)^n=x^{m \times n}\]

OpenStudy (michele_laino):

using the rule of power of power, we can rewrite your equation like this: \[\Large {\left( {{8^9}} \right)^p} = {8^{9p}} = {8^{18}}\]

OpenStudy (anonymous):

and 9+9 is 18!!

OpenStudy (anonymous):

oh

Nnesha (nnesha):

no when would you add the exponents ?

OpenStudy (anonymous):

sorry what?

Nnesha (nnesha):

when you multiply same bases you should `add` their exponents \[\huge\rm x^m \times x^n=x^{m + n}\] when you divide same bases yo should `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] \[\huge\rm (x^m)^n=x^{m \times n}\]

OpenStudy (anonymous):

its 2 then, right??

Nnesha (nnesha):

great!

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

thanks:DDDD

Nnesha (nnesha):

np :=)

OpenStudy (michele_laino):

so you have to solve this equation: \[\Large 9p = 18\]